OpenStudy (anonymous):

the legs of a right triangle mesured 6 in, and 8. in. what is the langth of the hypotenuse

OpenStudy (anonymous):

Use the Pythagorean theorem. a^2+b^2=c^2 c being the hypotenuse so 6^2+8^2=c^2 now solve

OpenStudy (anonymous):

so whats the anwer?

OpenStudy (anonymous):

A=6 B=8 C=10

OpenStudy (anonymous):

ok help with more

OpenStudy (anonymous):

one leg of a right triangle mesures 15 m, and the hypotenuse is 17 m, what is the length of the other leg?